# Nonlinearity-induced symmetry breaking in a system of two parametrically driven Kerr-Duffing oscillators

### v1

@@ V Broken-symmetry solutions of the symmetric case
Figure 4: An example of the behavior of the system in the presence of a broken-symmetry solution corresponding to the parameters shown in Fig. 3 (a), i.e. parameters: $g=1/5$, $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$. The stable solutions for the two amplitudes at different values of the detuning are shown. Before the bifurcation, we have the symmetric solutions $|\bar{z}_{0}|$, which are placed on a diagonal through the origin: the two modes can have either the same phase or a phase shift by $\pi$ (see text). After the bifurcation, we obtained different stable paired solutions. For example the solution $z_{1}=|\bar{z}_{h}|e^{i\varphi_{h,+}}$ in quadrant II can be paired either with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,+}}$ in quadrant I or with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,-}}$ in quadrant III.
@@
Figure 5: Phase diagram for the symmetric resonators with equal resonant frequency and drive strength $\delta\Omega_{T}=10$. The symmetric solutions $|\bar{z}_{1}|=|\bar{z}_{2}|$ are stable in the white region, whereas the broken symmetry solutions are stable in the green region.
@@
In Fig. 5, we show the complete phase diagram. The white areas represent regions where symmetric solutions are dynamically stable. In contrast, red areas indicate regions of dynamic instability for these solutions. Green areas denote the existence and dynamic stability of broken-symmetry solutions. When green overlaps with red, the broken-symmetry solution is the only stable configuration. Conversely, where green overlaps with white, both symmetric and broken-symmetry solutions remain dynamically stable. We found the stability regions for the broken-symmetry solutions numerically, whereas the symmetric solutions are obtained from analytical formulas that we analyze in detail in the following. In the white regions, we find that also other type solutions, for example, the solution with $z_{1}=0$ and $z_{2}\neq 0$, are still possible. A similar phase diagram occurs if we vary the symmetric drive strength $\delta\Omega_{T,1}=\delta\Omega_{T,1}$ where the critical coupling at $g_{c}\approx 1/3$ remains unchanged.
$z_{0}=\pm i\,\frac{\delta\Omega+\delta\Omega_{T}}{1+3g}$ (15)
$a$ $=2\left(1+g\right){\left|z_{0}\right|}^{2}-\delta\Omega\,,$ (16)
$b$ $=-\left(1+g\right){\left|z_{0}\right|}^{2}+\delta\Omega_{T}\,,$ (17)
$c$ =±$=\pm 4g{\left|z_{0}\right|}^{2}\,,$ (18)
$d$ $=\mp 2g{\left|z_{0}\right|}^{2}\,.$ (19)
with the sign $\pm$ ()minus-or-plus($\mp$)( ∓ ) depending on the phases of the paired solution, e.g. the same $\bar{z}_{1}=\bar{z}_{2}$ or different phase $\bar{z}_{1}=-\bar{z}_{2}$. Recalling that $\delta z_{n}=\delta Q_{n}-i\delta P_{n}$ and changing the variables $\bar{Q}=\delta Q_{1}+\delta Q_{2}$, $\Delta Q=\delta Q_{1}-\delta Q_{2}$, $\bar{P}=\delta P_{1}+\delta P_{2}$, $\Delta P=\delta P_{1}-\delta P_{2}$, one can obtain the expansion of the Hamiltonian around the symmetric stable solution as the same of two quadratic Hamiltonian
$H_{RWA}^{(2)}=H_{\Sigma}+H_{\Delta}\,.$ (20)
@@
$H_{\Sigma}$ $=\frac{1}{2}\Omega_{\bar{P}}\,\,\bar{P}^{2}+\frac{1}{2}\Omega_{\bar{Q}}\,\,\bar{Q}^{2}\,,$ (21)
$H_{\Delta}$ $=\frac{1}{2}\Omega_{\Delta P}\,\,\Delta P^{2}+\frac{1}{2}\Omega_{\Delta Q}\,\,\Delta Q^{2}\,.$ (22)
in which the conjugate variables are $\left(\bar{Q},\bar{P}\right)$ and $\left(\Delta Q,\Delta P\right)$ with the following dynamical equations
$\dot{\bar{P}}=\partial H_{\Sigma}/\partial\bar{Q}\,,\quad\dot{\bar{Q}}=-\partial H_{\Sigma}/\partial\bar{P}\,,$ (23)
$\Delta\dot{P}=\partial H_{\Delta}/\partial\Delta Q\,,\quad\Delta\dot{Q}=-\partial H_{\Delta}/\partial\Delta P\,.$ (24)
The system is dynamically stable as long as the product of the two coefficients in each quadratic Hamiltonian is positive, namely $\Omega_{\bar{P}}\Omega_{\bar{Q}}>0$ and $\Omega_{\Delta P}\Omega_{\Delta Q}>0$.
For example, for $\bar{z}_{1}=\bar{z}_{2}$, the coefficients are
$\Omega_{\bar{P}}$ $=2\left(\delta\Omega+\delta\Omega_{T}\right)\,,$ (25)
$\Omega_{\bar{Q}}$ $=2\delta\Omega_{T}\,,$ (26)
and
$\Omega_{\Delta P}$ $=2\,\frac{1-3g}{1+3g}\left(\delta\Omega+\delta\Omega_{T}\right)\,,$ (27)
$\Omega_{\Delta Q}$ $=2\,\frac{\delta\Omega_{T}}{1+3g}\left[1-g\left(2\frac{\delta\Omega}{\delta\Omega_{T}}-1\right)\right]\,.$ (28)
The quadratic Hamiltonian around the stationary point $\bar{z}_{1}=-\bar{z}_{2}$ has a similar form which can be written by exchanging the coefficient $\Omega_{\bar{P}}$ with $\Omega_{\Delta P}$ and the coefficient $\Omega_{\bar{Q}}$ with $\Omega_{\Delta Q}$.
VI Summary
In conclusion, we have studied the dynamics of a system of two parametrically driven nonlinear oscillators coupled by a nonlinear interaction. Several cases have been identified for which analytic solutions for the stationary state are available, while dynamic stability can be numerically assessed. As a most interesting result, in the symmetric case where the oscillators have equal parameters and are subject to equal driving forces, stationary solutions that spontaneously break the symmetry are found, leading to intriguing multistability phenomena. The natural next step will be to investigate the classical and quantum fluctuation properties around the critical points for this multistability phenomenon and the dependence of the switching rate between different solutions on the distance from the critical point.
Another interesting perspective is given by the recent advances in the field of superconducting circuits [32], where these types of non-linear oscillators can be implemented by using various Josephson junction-based technologies [33]. In this framework, the system can reach a regime where quantum effects become relevant. The pitchfork bifurcation leading to symmetry breaking in the steady-state solutions can eventually be described by a spontaneous symmetry-breaking dissipative phase transition in the quantum regime. Our findings can thus provide an interesting starting point to explore a new type of dissipative phase transition [34, 35, 36, 37, 38, 39] with possibly interesting non-classical states.
Acknowledgements.
The work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 25217212 - SFB 1432. G.R. and I.C. acknowledge financial support from the Provincia Autonoma di Trento, from the Q@TN initiative and from PNRR MUR project PE0000023-NQSTI.
Appendix A The model Hamiltonian
In this appendix we derive Eqs. (1),(2). We start from the model Hamiltonian that describes two nonlinear resonators of frequency $\omega_{n}$ with $n=1,2$ and Kerr-Duffing parameter $\kappa_{n}$ on which one applies one or two parametric external drives at frequency $2\omega_{d}$ and of strength $\mu_{n}$. The two modes also have a dispersive (or parametric) interaction of coupling strength $\lambda$. The Hamiltonian of the system reads
$H=\sum_{n=1}^{2}\left[\frac{p_{n}^{2}}{2}+\left(\frac{\omega_{n}^{2}}{2}+\mu_{n}\cos\left(2\omega_{d}t\right)\right)q_{n}^{2}+\frac{\kappa_{n}}{4}q_{n}^{4}\right]+\frac{\lambda}{2}q_{1}^{2}q_{2}^{2}$ (30)
@@ Appendix A The model Hamiltonian
Using the rotating wave approximation (RWA) with $\omega_{d}\gg|\omega_{d}-\omega_{i}|$, the fast oscillating components of the motion in the rotating frame can be neglected, and the following time-independent equation for the quadratures is obtained
$\frac{\partial X_{n}}{\partial t}=\frac{\partial H_{RWA}}{\partial Y_{n}}-\Gamma_{n}X_{n}\,,\quad\frac{\partial Y_{n}}{\partial t}=-\frac{\partial H_{RWA}}{\partial X_{n}}-\Gamma_{n}Y_{n}\,.$ (33)
The conservative dynamic is described by the effective Hamiltonian $H_{RWA}$. Before giving the explicit form of $H_{RWA}$, it is useful to scale the quadrature according to the following way
@@
where $Q_{n},P_{n}$ are dimensionless. Then, the effective Hamiltonian $H_{RWA}$ can be cast as
$\tilde{H}_{RWA}=\frac{3\sqrt{\gamma_{1}\gamma_{2}}}{8\omega_{d}\Gamma_{1}\Gamma_{2}}H_{RWA}=\alpha\,\tilde{H}_{1}+\frac{1}{\alpha}\,\tilde{H}_{2}+g\,\tilde{H}_{int}$ (35)
@@
$g$ represents the scaled coupling strength for the interaction, whereas $\alpha$ is the asymmetry parameter for the two modes. The dimensionless Hamiltonians are given by
@@
$\delta\Omega_{n}=\frac{\omega_{d}-\omega_{n}}{\Gamma_{n}}\,,\qquad\delta\Omega_{T,n}=\frac{2\mu_{n}}{4\omega_{d}\Gamma_{n}}$ (39)
@@
Hereafter, we assume the resonators are almost equal and the asymmetry factor is $\alpha\approx 1$. More precisely, we assume that the $\Gamma_{1}\approx\Gamma_{2}$ which implies $\kappa_{1}\approx\kappa_{2}$ for $\alpha=1$. This is valid for $\Gamma_{i}\gg|\Gamma_{1}-\Gamma_{2}|$ and $\kappa_{i}\gg|\kappa_{1}-\kappa_{2}|$, i.e. the difference between the parameters are small corrections. Therefore, to simplify the notation, we can set $\Gamma_{1}=\Gamma_{2}=\Gamma$ and $\kappa_{1}\approx\kappa_{2}=\kappa$ By scaling the time as $\tau=\Gamma t$ we obtain Eqs. (1),(2).
@@ Appendix B The noninteracting case $g=0$ and the steady-state trivial solution $\bar{z}_{1}=\bar{z}_{2}=0$
The solution $\bar{u}^{(0)}_{1,+}$ corresponds to the the resonator state when it starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is completely out of phase with respect to the drive. By analyzing the dynamics of the harmonic fluctuations around the stationary point, this solution is always dynamically stable above the threshold. The solution $\bar{u}^{(0)}_{1,-}$ appears above the detuning threshold $\delta\Omega>\delta\Omega_{T,1}$ and it is in phase respect to the drive: This solution is always dynamically unstable above the corresponding threshold. The trivial solution $\bar{u}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
@@ Appendix C Solutions with $\bar{z}_{2}=0$
The equations for the fluctuations of the two resonators are decoupled, and in particular, the dynamical equations for the fluctuations of the resonator with $(\bar{z}_{1}\neq 0$ remain unchanged respect to the noninteracting case $g=0$.
However, the two resonators are correlated as the finite value of the amplitude of the first resonator $\bar{z}_{1}$ affects the fluctuations $\delta z_{2}$ of the second one with $\bar{z}_{2}=0$. Therefore, the stability range can differ from that of the noninteracting case. As a consequence, the stationary pair solution, which is stable for the case $g=0$ $\left(\bar{z}_{1,+}=\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}},\bar{z}_{2}=0\right)$ can become unstable. For example, in the regime $\omega_{1}=\omega_{2}$, $\delta\Omega_{T,2}<g\delta\Omega_{T,1}$ and $g<1/2$, the zero amplitude solution for the second resonator $\bar{z}_{2}=0$ becomes unstable in the detuning range $\delta\Omega_{T,2}<\delta\Omega<\left(2g\delta\Omega_{T,1}-\delta\Omega_{T,2}\right)/(1-2g)$, beyond the range of the noninteracting case $|\delta\Omega|<\delta\Omega_{T,2}$.
@@ Appendix D Analytic solutions for $\delta\Omega_{T,1}>0$ and $\delta\Omega_{T,2}=0$
The solutions in table 1 are possible solutions. However, as second step, we analyze the dynamical stability of these stationary solutions. In general, the region of stability does not coincide with the detuning threshold at which the stationary solution appears, and we computed such ranges numerically.
In Fig. 1, we report an example to illustrate the typical behavior of the system with the discontinuity in the first derivative of $\bar{z}_{1}$ as a function of detuning and the parametric-like threshold for $\bar{z}_{2}$.

### v2

@@ VI Broken-symmetry solutions for symmetric drive
Figure 4: An example of the behavior of the system in the presence of a broken-symmetry solution corresponding to the parameters shown in Fig. 3 (a). Parameters: $g=1/5$, $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$. The stable solutions for the two amplitudes at different values of the detuning are shown. Before the bifurcation, we have the symmetric solutions $|\bar{z}_{0}|$, which are placed on a diagonal through the origin: the two modes can have either the same phase or a phase shift by $\pi$ (see text). After the bifurcation, we obtained different stable paired solutions. For example the solution $z_{1}=|\bar{z}_{h}|e^{i\varphi_{h,+}}$ in quadrant II can be paired either with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,+}}$ in quadrant I or with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,-}}$ in quadrant III.
@@
Figure 5: Phase diagram for the symmetric resonators with equal resonant frequency and drive strength $\delta\Omega_{T}=10$. The symmetric solutions $|\bar{z}_{1}|=|\bar{z}_{2}|$ are stable (unstable) in the white (red) region. The broken symmetry solutions are stable in the green region.
@@
In Fig. 5, we show the complete phase diagram. The white areas represent regions where symmetric solutions are dynamically stable. In contrast, red areas indicate regions of dynamic instability for these solutions. Green areas denote the existence and dynamic stability of broken-symmetry solutions. When green overlaps with red, the broken-symmetry solution is the only stable configuration. Conversely, where green overlaps with white, both symmetric and broken-symmetry solutions remain dynamically stable.
We found the stability regions for the broken-symmetry solutions numerically, whereas the symmetric solutions are obtained from analytical formulas discussed in the previous section. A similar phase diagram occurs if we vary the symmetric drive strength $\delta\Omega_{T,1}=\delta\Omega_{T,2}$ where the critical coupling at $g_{c}\approx 1/3$ remains unchanged.
Given the parametric form of the driving and of the coupling, one must not forget that solutions where at least one oscillator is at rest, $z_{1}=0$ or $z_{2}=0$, are always possible.
VII Summary
In conclusion, we have studied the dynamics of a system of two parametrically driven nonlinear oscillators coupled by a nonlinear interaction. Several cases have been identified for which analytic solutions for the stationary state are available, while dynamic stability can be numerically assessed. As a most interesting result, in the symmetric case where the oscillators have equal parameters and are subject to equal driving forces, stationary solutions that spontaneously break the symmetry are found, leading to intriguing multistability phenomena. The natural next step will be to investigate the fluctuation properties around the critical points for this multistability phenomenon and the dependence of the switching rate between different solutions on the distance from the critical point.
Another interesting perspective is given by the recent advances in the field of superconducting circuits [36], where these types of nonlinear oscillators can be implemented by using various Josephson junction-based technologies [1, 37]. In this framework, the system can reach a regime where quantum effects become relevant. The pitchfork bifurcation leading to symmetry breaking in the steady-state solutions can eventually be described by a spontaneous symmetry-breaking dissipative phase transition in the quantum regime where the switching process is triggered by quantum fluctuations. Our findings can thus provide an interesting starting point to explore new types of dissipative phase transition in which the broken symmetry phase is multistable [38, 39, 40, 41, 42, 43].
Acknowledgements.
The work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 25217212 - SFB 1432. G.R. and I.C. acknowledge financial support by the PE0000023-NQSTI project by the Italian Ministry of University and Research, co-funded by the European Union - NextGeneration EU, and from Provincia Autonoma di Trento (PAT), partly via the Q@TN initiative. D.D.B. acknowledges funding from the European Union - NextGeneration EU, ”Integrated infrastructure initiative in Photonic and Quantum Sciences” - I-PHOQS [IR0000016, ID D2B8D520, CUP B53C22001750006].
Appendix A The model Hamiltonian
In this appendix we derive Eqs. (1),(2). We start from the model Hamiltonian that describes two nonlinear resonators of frequency $\omega_{n}$ with $n=1,2$ and Kerr-Duffing parameter $\kappa_{n}$ on which one applies one or two parametric external drives at frequency $2\omega_{d}$ and of strength $\mu_{n}$.
The two modes also have a dispersive interaction of coupling strength $\lambda$. The Hamiltonian of the system reads
$H=\sum_{n=1}^{2}H_{n}+\frac{\lambda}{2}q_{1}^{2}q_{2}^{2}$ (30)
@@ Appendix A The model Hamiltonian
$H_{n}=\frac{p_{n}^{2}}{2}+\left(\frac{\omega_{n}^{2}}{2}+\mu_{n}\cos\left(2\omega_{d}t\right)\right)q_{n}^{2}+\frac{\kappa_{n}}{4}q_{n}^{4}$ (31)
@@
Using the rotating wave approximation (RWA) with $\omega_{d}\gg|\omega_{d}-\omega_{n}|$, the fast oscillating components of the motion in the rotating frame can be neglected, and the following time-independent equation for the quadratures is obtained
$\frac{\partial X_{n}}{\partial t}=\frac{\partial H_{\mathrm{RWA}}}{\partial Y_{n}}-\Gamma_{n}X_{n}\,,\quad\frac{\partial Y_{n}}{\partial t}=-\frac{\partial H_{\mathrm{RWA}}}{\partial X_{n}}-\Gamma_{n}Y_{n}\,.$ (34)
The conservative dynamic is described by the effective Hamiltonian $H_{\mathrm{RWA}}$. Before giving the explicit form of $H_{\mathrm{RWA}}$, it is useful to scale the quadrature according to the following way
@@
where $Q_{n},P_{n}$ are dimensionless. Then, the effective Hamiltonian $H_{\mathrm{RWA}}$ can be cast as
$\tilde{H}_{\mathrm{RWA}}=\frac{3\sqrt{\gamma_{1}\gamma_{2}}}{8\omega_{d}\Gamma_{1}\Gamma_{2}}H_{\mathrm{RWA}}=\alpha\,\tilde{H}_{1}+\frac{1}{\alpha}\,\tilde{H}_{2}+g\,\tilde{H}_{int}$ (36)
@@
$g$ represents the scaled coupling strength for the interaction, whereas $\alpha$ is the asymmetry parameter. The dimensionless Hamiltonians are given by
@@
$\delta\Omega_{n}=\frac{\omega_{d}-\omega_{n}}{\Gamma_{n}}\,,\qquad\delta\Omega_{T,n}=\frac{\mu_{n}}{2\omega_{d}\Gamma_{n}}$ (40)
@@
In the main text, we assumed that the resonators are almost equal and the asymmetry factor is $\alpha\approx 1$. More precisely, we consider $\Gamma_{1}\approx\Gamma_{2}$ which implies $\kappa_{1}\approx\kappa_{2}$ for $\alpha=1$. This is valid for $\Gamma_{n}\gg|\Gamma_{1}-\Gamma_{2}|$ and $\kappa_{n}\gg|\kappa_{1}-\kappa_{2}|$, i.e. the difference between the parameters are small corrections. Therefore, to simplify the notation, we set $\Gamma_{1}\approx\Gamma_{2}=\Gamma$ and $\kappa_{1}\approx\kappa_{2}=\kappa$. By scaling the time as $\tau=\Gamma t$ we obtain Eqs. (1),(2).
@@ Appendix B The noninteracting case $g=0$ and the steady-state trivial solution $\bar{z}_{1}=\bar{z}_{2}=0$
The solution $\bar{z}^{(0)}_{1,+}$ corresponds to the the resonator state when it starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is completely out of phase with respect to the drive. By analyzing the dynamics of the harmonic fluctuations around the stationary point, this solution is always dynamically stable above the threshold. The solution $\bar{z}^{(0)}_{1,-}$ appears above the detuning threshold $\delta\Omega>\delta\Omega_{T,1}$ and it is in phase respect to the drive: This solution is always dynamically unstable above the corresponding threshold. The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
@@ Appendix C Solutions with $\bar{z}_{2}=0$
The equations for the fluctuations of the two resonators are decoupled, and in particular, the dynamical equations for the fluctuations of the resonator with $\bar{z}_{1}\neq 0$ remain unchanged respect to the noninteracting case $g=0$.
However, the two resonators are correlated as the finite value of the amplitude of the first resonator $\bar{z}_{1}$ affects the fluctuations $\delta z_{2}$ of the second one with $\bar{z}_{2}=0$. Therefore, the stability range can differ from that of the noninteracting case. As a consequence, the stationary pair solution, which is stable for the case $g=0$ $\left(\bar{z}_{1,+}=\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}},\bar{z}_{2}=0\right)$ can become unstable. For example, in the regime $\omega_{1}=\omega_{2}$, $\delta\Omega_{T,2}<g\delta\Omega_{T,1}$ and $g<1/2$, the zero amplitude solution for the second resonator $\bar{z}_{2}=0$ becomes unstable in the detuning range $\delta\Omega_{T,2}<\delta\Omega<\left(2g\delta\Omega_{T,1}-\delta\Omega_{T,2}\right)/(1-2g)$, well beyond the range of the noninteracting case $|\delta\Omega|<\delta\Omega_{T,2}$.
@@ Appendix D Analytic solutions for $\delta\Omega_{T,1}>0$ and $\delta\Omega_{T,2}=0$
The solutions in table 1 are possible solutions. However, as second step, we analyze the dynamical stability of these stationary solutions. In general, the region of stability does not coincide with the detuning threshold at which the stationary solution appears, and we must hence compute the stability ranges numerically.
In Fig. 1, we reported an example to illustrate the typical behavior of the system with the discontinuity in the first derivative of $\bar{z}_{1}$ as a function of detuning and the parametric-like threshold for $\bar{z}_{2}$.